# Modified Marrone-Treanor dissociation model: formulation and verification for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL respectively

### v1

@@ 8 Conclusion
This paper has presented a new two-temperature model for air dissociation. Its outstanding feature is that all its rate data have been derived exclusively from fitting quasiclassical trajectory (QCT) calculations on highly accurate ab initio potential energy surfaces (PESs), generated by University of Minnesota computational chemists over the last decade. The reaction set considered in this derivation encompasses oxygen and nitrogen molecule dissociation with various collision partners, but the full model comprises additional rate data for reactions involving nitric oxide relevant to a 5-species air mixture.
@@
The analysis of QCT datasets for reactant molecule internal energy states sampled from Boltzmann distributions over wide $(T,T_{v})$-ranges revealed that vibrational energy change per dissociation, as well as vibration-level specific dissociation probability density (referred to as support factor) collapse to nearly unique curves when normalized by dissociation energy $D_{0}$ in a suitable way. The analysis also highlights that vibrational energy of the reactant molecules has a much greater influence on dissociation probability than that stored in rotation and collapses when plotted against “remaining energy”, i.e. the gap between the molecule’s potential energy barrier and its internal (vibrational+rotational) energy. The QCT results indicate that a rather simple two-temperature model is able to capture the major relevant physics of shock-heated dissociating diatomic species and that our proposed modified form of the classical the Marrone-Treanor model (called MMT) fits all QCT data (based on Boltzmann internal energy distributions) reasonably well. We find that simple correction factors can account for non-Boltzmann depletion effects observed in direct molecular simulations (DMS) employing the same PESs. Furthermore, the simple functional form proposed for these correction factors ensures that such depletion effects vanish as the mixture approaches chemical equilibrium.
@@
Next, we verified our CFD implementation of the MMT model against DMS reference solutions in space-homogeneous isothermal and adiabatic heat baths representative of post-shock conditions in hypersonic flows. Simulations of $\mathrm{N_{2}/N}$ and $\mathrm{O_{2}/O}$ mixtures show that, with our appropriately calibrated kinetic rate parameters, the two-temperature formulation of the MMT model closely matches DMS results even when dissociation takes place under vibrational nonequilibrium conditions. The MMT model’s mutually consistent expressions for nonequilibrium dissociation rate coefficient and associated vibrational energy removal term ensure that the correct amount of energy is transferred between the vibrational and trans-rotational modes at all times. This feature is particularly important for achieving close agreement between CFD-MMT and DMS under adiabatic conditions, where $T$ and $T_{\mathrm{v}}$ both vary in response to dissociation reactions and vibrational relaxation altering the gas state.
The resulting MMT model accurately reproduces the major features observed in direct molecular simulations, but remains computationally inexpensive enough to be employed in large-scale CFD simulations. An upcoming companion article fully parameterizes the MMT model for 5-species air, including vibrational relaxation rates and verifies its accuracy against DMS relaxation results in isothermal and adiabatic heat baths.

### v2

@@ 8 Conclusion
In this paper we have presented a new two-temperature model for air species dissociation. The reaction set considered in this derivation encompasses $\mathrm{N_{2}}$ dissociation with collision partners $\mathrm{M=\{N_{2},N,O_{2}\}}$ and $\mathrm{O_{2}}$ dissociation with partners $\mathrm{M=\{O_{2},O,N_{2}\}}$. The new model’s outstanding feature is that all its parameters were derived exclusively by fitting to rate data derived from QCT calculations on highly accurate ab initio potential energy surfaces. The entire PES set was provided by the Computational Chemistry group at the University of Minnesota.
Our analysis of QCT datasets for reactant molecule internal energy states sampled from Maxwell-Boltzmann distributions over wide $(T,T_{\mathrm{v}})$ ranges revealed that vibrational energy change per dissociation, as well as the vibration-level-specific dissociation probability density (referred to as support factor) collapse to nearly unique curves when normalized by dissociation energy $D_{0}$ in a suitable way. The analysis also highlighted that a reactant diatom’s vibrational energy has a much greater influence on its dissociation probability than that stored in its rotational motion.
@@
In aggregate, the QCT results suggest that a rather simple two-temperature model is able to capture the major relevant physics of shock-heated dissociating diatomic species and that our proposed modified form of the classical Marrone-Treanor model (abbreviated as MMT) fits all the QCT data based on Boltzmann internal energy distributions reasonably well. We find that simple correction factors can account for non-Boltzmann depletion effects observed in DMS, or state-resolved master equation calculations employing the same ab initio PESs. Furthermore, a simple functional form proposed for these correction factors ensures that such depletion effects vanish as the mixture approaches chemical equilibrium.
@@
We benchmarked our CFD implementation of the MMT model against DMS reference solutions in space-homogeneous isothermal and adiabatic heat baths representative of post-shock conditions in hypersonic flows. Simulations of $\mathrm{N_{2}/N}$ and $\mathrm{O_{2}/O}$ mixtures show that, with our appropriately calibrated kinetic rate parameters, the two-temperature formulation of the MMT model closely matches DMS results even when dissociation takes place under vibrational nonequilibrium conditions. The MMT model’s mutually consistent expressions for nonequilibrium dissociation rate coefficient and associated vibrational energy removal term ensure that the correct amount of energy is transferred between the vibrational and trans-rotational modes at all times. This feature is particularly important for achieving close agreement between CFD-MMT and DMS predictions under adiabatic conditions, where $T$ and $T_{\mathrm{v}}$ both vary dynamically in response to dissociation coupled to vibrational relaxation.
The proposed MMT model accurately reproduces the major features observed in DMS calculations, but remains computationally inexpensive enough to be employed in large-scale CFD simulations. In a follow-up article we will parameterize the MMT model for 5-species air, including new QCT data for the Zeldovich exchange reactions, and benchmark its predictions against DMS reference results for partially dissociated air.